# Adaptive Pruning of Pretrained Transformer via Differential Inclusions

### v1

@@ Abstract
Large transformers have demonstrated remarkable success, making it necessary to compress these models to reduce inference costs while preserving their performance. Current compression algorithms prune transformers at fixed compression ratios, requiring a unique pruning process for each ratio, which results in high computational costs. In contrast, we propose pruning of pretrained transformers at any desired ratio within a single pruning stage, based on a differential inclusion for a mask parameter. This dynamic can generate the whole regularization solution path of the mask parameter, whose support set identifies the network structure. Therefore, the solution path identifies a Transformer weight family with various sparsity levels, offering greater flexibility and customization. In this paper, we introduce such an effective pruning method, termed SPP (Solution Path Pruning). To achieve effective pruning, we segment the transformers into paired modules, including query-key pairs, value-projection pairs, and sequential linear layers, and apply low-rank compression to these pairs, maintaining the output structure while enabling structural compression within the inner states. Extensive experiments conducted on various well-known transformer backbones have demonstrated the efficacy of SPP.
@@ 1 Introduction
Transformers have succeeded in various tasks due to their scalability, parallel processing abilities, and capacity to learn complex data patterns. Pretrained transformers, in particular, perform well on downstream tasks by leveraging large datasets during training. These models are highly effective for transfer learning, allowing fine-tuning for different tasks and delivering strong performance in many natural language processing applications. However, their large size, often with billions of parameters, makes them difficult to deploy on low-cost hardware. Despite their widespread use and impressive performance Radford et al. (2021); Touvron et al. (2021a), running transformers on lightweight devices like phones and laptops remains challenging due to computational constraints. Therefore, compressing transformers to run efficiently on affordable hardware is crucial.
Various techniques have been developed to compress transformers while preserving their performance. These include weight sharing Lan et al. (2019), low-rank factorization Yu et al. (2017), quantization Gong et al. (2014); Polino et al. (2018); Tao et al. (2022), knowledge distillation Hinton et al. (2015); Yuan et al. (2019); Touvron et al. (2020); Liu et al. (2020), and pruning Yu & Xiang (2023); Yang et al. (2023); Shi et al. (2023); Yin et al. (2023). Pruning methods typically involve structured pruning, wherein entire neurons or attention heads are removed based on importance, followed by fine-tuning to regain accuracy. Despite the array of pruning methods developed for transformers, current algorithms face a fundamental challenge: they are tailored to achieve a predefined pruning ratio by the end of the pruning phase.
@@
Our exploration of transformer structures adopts a novel and efficient approach, automatically learning compression structures throughout training, thus enjoying greater flexibility and customization. Currently, no other algorithm can theoretically guarantee optimization convergence while producing sparse models with different levels of sparsity during training. For example, the Upop Shi et al. (2023) algorithm only makes the model structure sparse at the final step, with the mask becoming zero only at that point. The OFA Cai et al. (2019) method lacks a convergence guarantee and simply fixes the mask updates during training to achieve a sparse network. SPP  addresses this by using the solution path to obtain sparse structures without affecting convergence, providing a stronger theoretical foundation. Notably, our method can be extended to prune large language models (LLMs) with one-shot post training.
In this paper, we applied SPP to the classification task dataset ImageNet Deng et al. (2009) using the DeiT Touvron et al. (2021a) backbone. Furthermore, we also extended the method to image and text retrieval datasets, using CLIP models Radford et al. (2021). Our contributions are summarized as follows:
  1. 1.
@@ 2 Related work
Over recent years, various structured pruning methods have been developed. ViT-Slim Chavan et al. (2022) employs a single-shot training approach to search for optimal sub-structures in vision transformers. SAViT Chen et al. (2021) collaboratively prunes all components of Vision Transformers, integrating structure-aware interactions and using a Taylor-based optimization function for joint importance. UP-ViTs Yu & Wu (2023) introduces a unified pruning framework for vision transformers, focusing on comprehensive pruning of all components of ViT models and their variants.
The process of pruning a Transformer model is twofold: first targeting the MLP and then the Attention mechanism. Approaches such as WDpruning Yu et al. (2022) employ a mask-based technique. Specifically, a mask $M$ is defined to correspond to each column of the MLP’s weight matrix, and pruning is conducted by considering the gradient magnitudes of this mask. However, due to the complicated structure of the Attention mechanism, such strategies may fail.
WDpruning Yu et al. (2022) extends its paradigm by incorporating head-level pruning within the multi-head attention framework. Concurrently, methodologies like SAVIT and Upop advocate for the pruning of the input projection matrices, retaining the structural integrity of the query, key, and value matrices. Those approaches, however, lack flexibility and expandability. Because the matrices for the query, key, and value play distinct roles during the forward pass of the attention process.
@@
Mirror Descent Algorithm and Bregman Inverse Scale Space. Mirror Descent Algorithm (MDA) was first proposed by Nemirovskij & Yudin (1983) to solve constrained convex optimization, and can be seen as a generalized projected gradient descent Beck & Teboulle (2003) using Bregman distance $B_{\Omega}(u,v):=\Omega(u)-\Omega(v)-\langle\nabla\Omega(v),u-v\rangle$ induced by a convex and differentiable function $\Omega(.)$.
Convergence analysis for convex loss has been extended to stochastic versions Ghadimi & Lan (2013) and Nesterov acceleration Krichene et al. (2015); Su et al. (2016). For non-convex loss in deep learning, convergence to global optima for overparameterized networks has been established Azizan et al. (2019). For non-differentiable penalties, such as $\ell_{1}$ penalty for sparse recovery, Osher et al. (2016a); So et al. (2008) proposed Linearized Bregman Iteration (LBI), follows a discretized solution path of differential inclusions called Bregman Inverse Scale Space (ISS). These solution paths enjoy the inverse scale space property, which means important features such as the signal, will be learned earlier than non-important ones such as noise.
Mask-based pruning for pretrained CNN.Fu et al. (2020; 2022); Bungert et al. (2022) applied LBI to forwardly train a network from scratch, based on the lottery ticket hypothesis. By incorporating several training techniques tailored to the network architecture, the sparse network achieved comparable results to the fully dense network. In contrast, we propose a new differential inclusion for mask-based pruning, which can adaptively prune a pre-trained transformer. This method generates an iterative solution path that uncovers key sparse architectures early during training. Our method can perform consistently well across various backbones and datasets. Unlike ADMM Wahlberg et al. (2012); Boyd et al. (2011), which focuses on convergence, our differential inclusion dynamics aim at generating a whole solution path with various compression ratios.
@@ 3.2 Differential inclusion for regularization weight family
equation 8 is a special form of a mirror descent flow with $\Omega(\Gamma)$. Unlike previous works for mirror descent primarily concerned with convergence analysis, our main objective is to obtain a regularization solution path, where each solution along this solution corresponds to a sparse mask that defines the network’s sparse structure.
The differential inclusion in equation 8 generates a solution path for $(M_{t},\Gamma_{t})$, where $M_{t}$ follows a gradient descent flow to learn the activation values for the pre-trained weights $W_{0}$, while $\Gamma_{t}$ is updated via a mirror descent flow with the penalty $\Omega(\Gamma)$ to explore the sparse structure. During the updating process of $\Gamma_{t}$, important weights are learned earlier than non-important ones. Essentially, this is known as the inverse scale space property in inverse problems Burger et al. (2006); Osher et al. (2016b).
Similarly, starting from $\Gamma_{0}=\mathbf{0}$, $\Gamma_{t}$ identifies a family of network structures, where important weights will be learned in earlier epochs. Specifically, note that for each element $i$, $|V(i)|\leq 1$, if $\Gamma(i)=0$, and is equal to $1$ when it becomes non-zeros. Therefore, driven by the gradient descent in equation 8b that gives a dense solution $\lim_{t\to\infty}(M_{t},\Gamma_{t})=\arg\min_{M,\Gamma}\mathcal{L}_{\rho}(M,\Gamma)$, $V_{t}$ that begins from $V_{0}=\mathbf{0}$, will have more of its elements reaching the boundary of 1, making corresponding elements in $\Gamma_{t}$ becoming non-zeros. Compared to directly solving equation 7, our dynamics is more efficient in generating a network family with various sparsity levels and, hence is more flexible for deployment.
@@
The differential inclusion has been previously explored to forwardly train a sparse network from scratch Fu et al. (2020; 2022); Bungert et al. (2022), incorporating several deep learning techniques tailored to the multilayer perceptron and the convolutional neural network. However, these techniques may not be applicable to the Transformer, and thus may lead to suboptimal performance. In contrast, our dynamics involves backward pruning of a pre-trained network. Furthermore, instead of weight-based pruning, we employ mask-based pruning, which is significantly easier to train.
@@
where $\alpha$ is the step size and should be small enough to well approximate the original dynamics in  equation 8. In  equation 10b, the proximal operator $\mathrm{Prox}_{\Omega}(\cdot)$ is defined as:
@@
which gives $\Gamma(i)=1$ if $V(i)>2$, $=V(i)-1$ if $1<V\leq 2$, $=0$ otherwise. Running  equation 10 will obtain a solution path of $\Gamma_{k}$ with increasing levels, that define a family of weights dubbed as Transformer weight family:
@@
Model Method Top-1(%) Top-5(%) FLOPS(B) Params(M)
$\text{S}^{\mathrm{2}}$ ViTE-SmallChen et al. (2021) 79.2 - - $14.6_{{\color[rgb]{1,0,0}66.1\%}}$
GOHSPYin et al. (2023) 80.0 - $3.0_{{\color[rgb]{1,0,0}65.2\%}}$ $14.4_{{\color[rgb]{1,0,0}65.2\%}}$
PS-ViT-STang et al. (2022) 79.4 - $2.7_{{\color[rgb]{1,0,0}58.7\%}}$ $22.0_{{\color[rgb]{1,0,0}99.5\%}}$
ViTAS - E Su et al. (2022) 77.4 93.8 $2.7_{{\color[rgb]{1,0,0}58.7\%}}$ $12.6_{{\color[rgb]{1,0,0}57.0\%}}$
UpopShi et al. (2023) 79.6 94.8 $2.8_{{\color[rgb]{1,0,0}60.9\%}}$ $13.5_{{\color[rgb]{1,0,0}61.1\%}}$
UpopShi et al. (2023) 80.2 95.1 $3.2_{{\color[rgb]{1,0,0}69.6\%}}$ $15.7_{{\color[rgb]{1,0,0}71.0\%}}$
ViT-SlimChavan et al. (2022) 80.0 95.1 $3.3_{{\color[rgb]{1,0,0}71.7\%}}$ $15.7_{{\color[rgb]{1,0,0}71.0\%}}$
WDPruningYu et al. (2022) 78.6 94.4 $3.1_{{\color[rgb]{1,0,0}67.4\%}}$ $15.0_{{\color[rgb]{1,0,0}67.9\%}}$
X-PrunerYu & Xiang (2023) 78.9 94.2 $2.4_{{\color[rgb]{1,0,0}52.2\%}}$
OPTINKhaki & Plataniotis (2024) 79.2 - $3.2_{{\color[rgb]{1,0,0}68.4\%}}$
DeiT-Tiny Uncompressed 72.2 91.1 1.3 5.7
GOHSPYin et al. (2023) 70.2 - $0.9_{{\color[rgb]{1,0,0}69.2\%}}$ $4.0_{{\color[rgb]{1,0,0}70.2\%}}$
$\text{S}^{\mathrm{2}}$iTEChen et al. (2021) 70.1 - $1.0_{{\color[rgb]{1,0,0}76.9\%}}$ $4.2_{{\color[rgb]{1,0,0}73.7\%}}$
WDPruningYu et al. (2022) 70.3 89.8 $0.7_{{\color[rgb]{1,0,0}53.8\%}}$
PoWERGoyal et al. (2020) 69.4 89.2 $0.8_{{\color[rgb]{1,0,0}61.5\%}}$
UPDPLiu et al. (2024) 70.3 - $0.9_{{\color[rgb]{1,0,0}70.3\%}}$ $3.8_{{\color[rgb]{1,0,0}66.7\%}}$
MCFBai et al. (2023) 71.5 - $0.7_{{\color[rgb]{1,0,0}53.8\%}}$ $3.9_{{\color[rgb]{1,0,0}68.4\%}}$
OPTIN 71.3 - $0.9_{{\color[rgb]{1,0,0}70.3\%}}$
SCOPTang et al. (2020) 79.7 94.5 $10.2_{{\color[rgb]{1,0,0}58.3\%}}$ $58.3_{{\color[rgb]{1,0,0}67.3\%}}$
IA-RED Pan et al. (2021) 80.3 - $11.8_{{\color[rgb]{1,0,0}67.4\%}}$ $67.0_{{\color[rgb]{1,0,0}77.4\%}}$
PoWERGoyal et al. (2020) 80.1 94.6 $10.4_{{\color[rgb]{1,0,0}59.4\%}}$
X-Pruner 81.02 95.38 $8.5_{{\color[rgb]{1,0,0}48.6\%}}$
Table 1: The results of DeiTTouvron et al. (2021b) models.We compared the results with the other advanced methods.The reduce is the reduce of parameters.
@@ 4 Convergence
We present a theorem that guarantees the global convergence of our method, i.e. from any initialization, the sequence converges to a critical point of $\bar{\cal L}$. Our treatment extends the Block Coordinate Descent (BCD) studied in Zeng et al. (2019), with a crucial difference being the mirror descent involved in our method. Instead of the splitting loss in BCD, a new Lyapunov function is developed here to meet the Kurdyka-Łojasiewicz property Łojasiewicz (1963). Let $P:=(M,\Gamma)$. Following Huang & Yao (2018), our method algorithm can be rewritten as the following standard linearized Bregman iteration,
$P_{k+1}=\arg\min_{P}\left\{\langle P-P_{k},\alpha\nabla\bar{\cal L}(P_{k})\rangle+B_{\Psi}^{p_{k}}(P,P_{k})\right\}\vspace{-0.1in}$ (13)
@@
$\vspace{-0.1in}\Psi(P)$ $=\Omega_{\lambda}(\Gamma)+\frac{1}{2\kappa}\|P\|_{2}^{2}$
$=\Omega_{\lambda}(\Gamma)+\frac{1}{2\kappa}\|W\|_{2}^{2}+\frac{1}{2\kappa}\|\Gamma\|_{2}^{2},$ (14)
@@
$0<\alpha_{k}=\alpha<\frac{2}{\kappa(Lip*C+\nu^{-1})},\vspace{-0.1in}$
@@ 5 Experiments
GOHSPYin et al. (2023) 70.2 - $0.9_{{\color[rgb]{1,0,0}69.2\%}}$ $4.0_{{\color[rgb]{1,0,0}70.2\%}}$
$\text{S}^{\mathrm{2}}$iTEChen et al. (2021) 70.1 - $1.0_{{\color[rgb]{1,0,0}76.9\%}}$ $4.2_{{\color[rgb]{1,0,0}73.7\%}}$
WDPruningYu et al. (2022) 70.3 89.8 $0.7_{{\color[rgb]{1,0,0}53.8\%}}$
PoWERGoyal et al. (2020) 69.4 89.2 $0.8_{{\color[rgb]{1,0,0}61.5\%}}$
@@
ViT-SlimChavan et al. (2022) 98.7 - $3.3_{{\color[rgb]{1,0,0}71.7\%}}$ $15.6_{{\color[rgb]{1,0,0}70.6\%}}$
WDPruningYu et al. (2022) 98.1 - $2.8_{{\color[rgb]{1,0,0}60.9\%}}$ $14.9_{{\color[rgb]{1,0,0}67.4\%}}$

### v2

@@ Abstract
Large transformers have demonstrated remarkable success, making it necessary to compress these models to reduce inference costs while preserving their performance. Current compression algorithms prune transformers at fixed compression ratios, requiring a unique pruning process for each ratio, which results in high computational costs. In contrast, we propose pruning of pretrained transformers at any desired ratio within a single pruning stage, based on a differential inclusion for a mask parameter. This dynamic can generate the whole regularization solution path of the mask parameter, whose support set identifies the network structure. Therefore, the solution path identifies a Transformer weight family with various sparsity levels, offering greater flexibility and customization. In this paper, we introduce such an effective pruning method, termed SPP (Solution Path Pruning). To achieve effective pruning, we segment the transformers into paired modules, including query-key pairs, value-projection pairs, and sequential linear layers, and apply low-rank compression to these pairs, maintaining the output structure while enabling structural compression within the inner states. Extensive experiments conducted on various well-known transformer backbones have demonstrated the efficacy of SPP. Our code is available at https://github.com/yizhuoDi/Solution-Path-Pruning.
[…]
@@ 1 Introduction
Transformers have succeeded in various tasks due to their scalability, parallel processing abilities, and capacity to learn complex data patterns. Pretrained transformers, in particular, perform well on downstream tasks by leveraging large datasets during training. These models are highly effective for transfer learning, allowing fine-tuning for different tasks and delivering strong performance in many natural language processing applications. However, their large size, often with billions of parameters, makes them difficult to deploy on low-cost hardware. Despite their widespread use and impressive performance (Radford et al., 2021; Touvron et al., 2021a), running transformers on lightweight devices like phones and laptops remains challenging due to computational constraints. Therefore, compressing transformers to run efficiently on affordable hardware is crucial.
Various techniques have been developed to compress transformers while preserving their performance. These include weight sharing (Lan et al., 2019), low-rank factorization (Yu et al., 2017), quantization (Gong et al., 2014; Polino et al., 2018; Tao et al., 2022), knowledge distillation (Hinton et al., 2015; Yuan et al., 2019; Touvron et al., 2020; Liu et al., 2020), and pruning (Yu & Xiang, 2023; Yang et al., 2023; Shi et al., 2023; Yin et al., 2023). Pruning methods typically involve structured pruning, wherein entire neurons or attention heads are removed based on importance, followed by fine-tuning to regain accuracy. Despite the array of pruning methods developed for transformers, current algorithms face a fundamental challenge: they are tailored to achieve a predefined pruning ratio by the end of the pruning phase.
@@
Our exploration of transformer structures adopts a novel and efficient approach, automatically learning compression structures throughout training, thus enjoying greater flexibility and customization. Currently, no other algorithm can theoretically guarantee optimization convergence while producing sparse models with different levels of sparsity during training. For example, the Upop (Shi et al., 2023) algorithm only makes the model structure sparse at the final step, with the mask becoming zero only at that point. The OFA (Cai et al., 2019) method lacks a convergence guarantee and simply fixes the mask updates during training to achieve a sparse network. SPP  addresses this by using the solution path to obtain sparse structures without affecting convergence, providing a stronger theoretical foundation. Notably, our method can be extended to prune large language models (LLMs) with one-shot post training.
In this paper, we applied SPP to the classification task dataset ImageNet (Deng et al., 2009) using the DeiT (Touvron et al., 2021a) backbone. Furthermore, we also extended the method to image and text retrieval datasets, using CLIP models (Radford et al., 2021). Our contributions are summarized as follows:
@@ 2 Related work
Over recent years, various structured pruning methods have been developed. ViT-Slim (Chavan et al., 2022) employs a single-shot training approach to search for optimal sub-structures in vision transformers. SAViT (Chen et al., 2021) collaboratively prunes all components of Vision Transformers, integrating structure-aware interactions and using a Taylor-based optimization function for joint importance. UP-ViTs (Yu & Wu, 2023) introduces a unified pruning framework for vision transformers, focusing on comprehensive pruning of all components of ViT models and their variants.
The process of pruning a Transformer model is twofold: first targeting the MLP and then the Attention mechanism. Approaches such as WDpruning (Yu et al., 2022) employ a mask-based technique. Specifically, a mask $M$ is defined to correspond to each column of the MLP’s weight matrix, and pruning is conducted by considering the gradient magnitudes of this mask. However, due to the complicated structure of the Attention mechanism, such strategies may fail.
WDpruning (Yu et al., 2022) extends its paradigm by incorporating head-level pruning within the multi-head attention framework. Concurrently, methodologies like SAVIT and Upop advocate for the pruning of the input projection matrices, retaining the structural integrity of the query, key, and value matrices. Those approaches, however, lack flexibility and expandability. Because the matrices for the query, key, and value play distinct roles during the forward pass of the attention process.
@@
Mirror Descent Algorithm and Bregman Inverse Scale Space. Mirror Descent Algorithm (MDA) was first proposed by Nemirovskij & Yudin (1983) to solve constrained convex optimization, and can be seen as a generalized projected gradient descent (Beck & Teboulle, 2003) using Bregman distance $B_{\Omega}(u,v):=\Omega(u)-\Omega(v)-\langle\nabla\Omega(v),u-v\rangle$ induced by a convex and differentiable function $\Omega(.)$.
Convergence analysis for convex loss has been extended to stochastic versions (Ghadimi & Lan, 2013) and Nesterov acceleration (Krichene et al., 2015; Su et al., 2016). For non-convex loss in deep learning, convergence to global optima for overparameterized networks has been established (Azizan et al., 2019). For non-differentiable penalties, such as $\ell_{1}$ penalty for sparse recovery, Osher et al. (2016a); So et al. (2008) proposed Linearized Bregman Iteration (LBI), follows a discretized solution path of differential inclusions called Bregman Inverse Scale Space (ISS). These solution paths enjoy the inverse scale space property, which means important features such as the signal, will be learned earlier than non-important ones such as noise.
Mask-based pruning for pretrained CNN.Fu et al. (2020; 2022); Bungert et al. (2022) applied LBI to forwardly train a network from scratch, based on the lottery ticket hypothesis. By incorporating several training techniques tailored to the network architecture, the sparse network achieved comparable results to the fully dense network. In contrast, we propose a new differential inclusion for mask-based pruning, which can adaptively prune a pre-trained transformer. This method generates an iterative solution path that uncovers key sparse architectures early during training. Our method can perform consistently well across various backbones and datasets. Unlike ADMM (Wahlberg et al., 2012; Boyd et al., 2011), which focuses on convergence, our differential inclusion dynamics aim at generating a whole solution path with various compression ratios.
@@ 3.2 Differential inclusion for regularization weight family
Eq. 8 is a special form of a mirror descent flow with $\Omega(\Gamma)$. Unlike previous works for mirror descent primarily concerned with convergence analysis, our main objective is to obtain a regularization solution path, where each solution along this solution corresponds to a sparse mask that defines the network’s sparse structure.
The differential inclusion in Eq. 8 generates a solution path for $(M_{t},\Gamma_{t})$, where $M_{t}$ follows a gradient descent flow to learn the activation values for the pre-trained weights $W_{0}$, while $\Gamma_{t}$ is updated via a mirror descent flow with the penalty $\Omega(\Gamma)$ to explore the sparse structure. During the updating process of $\Gamma_{t}$, important weights are learned earlier than non-important ones. Essentially, this is known as the inverse scale space property in inverse problems (Burger et al., 2006; Osher et al., 2016b).
Similarly, starting from $\Gamma_{0}=\mathbf{0}$, $\Gamma_{t}$ identifies a family of network structures, where important weights will be learned in earlier epochs. Specifically, note that for each element $i$, $|V(i)|\leq 1$, if $\Gamma(i)=0$, and is equal to $1$ when it becomes non-zeros. Therefore, driven by the gradient descent in Eq. 8b that gives a dense solution $\lim_{t\to\infty}(M_{t},\Gamma_{t})=\arg\min_{M,\Gamma}\mathcal{L}_{\rho}(M,\Gamma)$, $V_{t}$ that begins from $V_{0}=\mathbf{0}$, will have more of its elements reaching the boundary of 1, making corresponding elements in $\Gamma_{t}$ becoming non-zeros. Compared to directly solving Eq. 7, our dynamics is more efficient in generating a network family with various sparsity levels and, hence is more flexible for deployment.
@@
The differential inclusion has been previously explored to forwardly train a sparse network from scratch (Fu et al., 2020; 2022; Bungert et al., 2022), incorporating several deep learning techniques tailored to the multilayer perceptron and the convolutional neural network. However, these techniques may not be applicable to the Transformer, and thus may lead to suboptimal performance. In contrast, our dynamics involves backward pruning of a pre-trained network. Furthermore, instead of weight-based pruning, we employ mask-based pruning, which is significantly easier to train.
@@
where $\alpha$ is the step size and should be small enough to well approximate the original dynamics in  Eq. 8. In  Eq. 10b, the proximal operator $\mathrm{Prox}_{\Omega}(\cdot)$ is defined as:
@@
which gives $\Gamma(i)=1$ if $V(i)>2$, $=V(i)-1$ if $1<V\leq 2$, $=0$ otherwise. Running  Eq. 10 will obtain a solution path of $\Gamma_{k}$ with increasing levels, that define a family of weights dubbed as Transformer weight family:
@@ 4 Convergence
We present a theorem that guarantees the global convergence of our method, i.e. from any initialization, the sequence converges to a critical point of $\bar{\cal L}$. Our treatment extends the Block Coordinate Descent (BCD) studied in Zeng et al. (2019), with a crucial difference being the mirror descent involved in our method. Instead of the splitting loss in BCD, a new Lyapunov function is developed here to meet the Kurdyka-Łojasiewicz property (Łojasiewicz, 1963). Let $P:=(M,\Gamma)$. Following Huang & Yao (2018), our method algorithm can be rewritten as the following standard linearized Bregman iteration,
$P_{k+1}=\arg\min_{P}\left\{\langle P-P_{k},\alpha\nabla\bar{\cal L}(P_{k})\rangle+B_{\Psi}^{p_{k}}(P,P_{k})\right\}$ (13)
@@
$\Psi(P)=\Omega_{\lambda}(\Gamma)+\frac{1}{2\kappa}\|P\|_{2}^{2}=\Omega_{\lambda}(\Gamma)+\frac{1}{2\kappa}\|W\|_{2}^{2}+\frac{1}{2\kappa}\|\Gamma\|_{2}^{2},$
@@
$0<\alpha_{k}=\alpha<\frac{2}{\kappa(Lip*C+\nu^{-1})},$ (16)
@@ 5 Experiments
Table 1: The results of DeiT(Touvron et al., 2021b) models.We compared the results with the other advanced methods.The reduce is the reduce of parameters.
Model Method Top-1(%) Top-5(%) FLOPS(B) Params(M)
$\text{S}^{\mathrm{2}}$ ViTE-Small(Chen et al., 2021) 79.2 - - $14.6_{{\color[rgb]{1,0,0}66.1\%}}$
GOHSP(Yin et al., 2023) 80.0 - $3.0_{{\color[rgb]{1,0,0}65.2\%}}$ $14.4_{{\color[rgb]{1,0,0}65.2\%}}$
PS-ViT-S(Tang et al., 2022) 79.4 - $2.7_{{\color[rgb]{1,0,0}58.7\%}}$ $22.0_{{\color[rgb]{1,0,0}99.5\%}}$
ViTAS - E (Su et al., 2022) 77.4 93.8 $2.7_{{\color[rgb]{1,0,0}58.7\%}}$ $12.6_{{\color[rgb]{1,0,0}57.0\%}}$
Upop(Shi et al., 2023) 79.6 94.8 $2.8_{{\color[rgb]{1,0,0}60.9\%}}$ $13.5_{{\color[rgb]{1,0,0}61.1\%}}$
Upop(Shi et al., 2023) 80.2 95.1 $3.2_{{\color[rgb]{1,0,0}69.6\%}}$ $15.7_{{\color[rgb]{1,0,0}71.0\%}}$
ViT-Slim(Chavan et al., 2022) 80.0 95.1 $3.3_{{\color[rgb]{1,0,0}71.7\%}}$ $15.7_{{\color[rgb]{1,0,0}71.0\%}}$
WDPruning(Yu et al., 2022) 78.6 94.4 $3.1_{{\color[rgb]{1,0,0}67.4\%}}$ $15.0_{{\color[rgb]{1,0,0}67.9\%}}$
X-Pruner(Yu & Xiang, 2023) 78.9 94.2 $2.4_{{\color[rgb]{1,0,0}52.2\%}}$
OPTIN(Khaki & Plataniotis, 2024) 79.2 - $3.2_{{\color[rgb]{1,0,0}68.4\%}}$
DeiT-Tiny Uncompressed 72.2 91.1 1.3 5.7
GOHSP(Yin et al., 2023) 70.2 - $0.9_{{\color[rgb]{1,0,0}69.2\%}}$ $4.0_{{\color[rgb]{1,0,0}70.2\%}}$
$\text{S}^{\mathrm{2}}$iTE(Chen et al., 2021) 70.1 - $1.0_{{\color[rgb]{1,0,0}76.9\%}}$ $4.2_{{\color[rgb]{1,0,0}73.7\%}}$
WDPruning(Yu et al., 2022) 70.3 89.8 $0.7_{{\color[rgb]{1,0,0}53.8\%}}$
PoWER(Goyal et al., 2020) 69.4 89.2 $0.8_{{\color[rgb]{1,0,0}61.5\%}}$
UPDP(Liu et al., 2024) 70.3 - $0.9_{{\color[rgb]{1,0,0}70.3\%}}$ $3.8_{{\color[rgb]{1,0,0}66.7\%}}$
MCF(Bai et al., 2023) 71.5 - $0.7_{{\color[rgb]{1,0,0}53.8\%}}$ $3.9_{{\color[rgb]{1,0,0}68.4\%}}$
OPTIN 71.3 - $0.9_{{\color[rgb]{1,0,0}70.3\%}}$
SCOP(Tang et al., 2020) 79.7 94.5 $10.2_{{\color[rgb]{1,0,0}58.3\%}}$ $58.3_{{\color[rgb]{1,0,0}67.3\%}}$
IA-RED (Pan et al., 2021) 80.3 - $11.8_{{\color[rgb]{1,0,0}67.4\%}}$ $67.0_{{\color[rgb]{1,0,0}77.4\%}}$
PoWER(Goyal et al., 2020) 80.1 94.6 $10.4_{{\color[rgb]{1,0,0}59.4\%}}$
X-Pruner 81.02 95.38 $8.5_{{\color[rgb]{1,0,0}48.6\%}}$
@@
GOHSP(Yin et al., 2023) 70.2 - $0.9_{{\color[rgb]{1,0,0}69.2\%}}$ $4.0_{{\color[rgb]{1,0,0}70.2\%}}$
$\text{S}^{\mathrm{2}}$iTE(Chen et al., 2021) 70.1 - $1.0_{{\color[rgb]{1,0,0}76.9\%}}$ $4.2_{{\color[rgb]{1,0,0}73.7\%}}$
WDPruning(Yu et al., 2022) 70.3 89.8 $0.7_{{\color[rgb]{1,0,0}53.8\%}}$
PoWER(Goyal et al., 2020) 69.4 89.2 $0.8_{{\color[rgb]{1,0,0}61.5\%}}$
@@
ViT-Slim(Chavan et al., 2022) 98.7 - $3.3_{{\color[rgb]{1,0,0}71.7\%}}$ $15.6_{{\color[rgb]{1,0,0}70.6\%}}$
WDPruning(Yu et al., 2022) 98.1 - $2.8_{{\color[rgb]{1,0,0}60.9\%}}$ $14.9_{{\color[rgb]{1,0,0}67.4\%}}$